OpenStudy (samigupta8):

The value of product of Re[Cis(2n-1)°] from n=1 to 45 Whole upon the product of Im[cis2(2n-1)°] from n=1 to n=45

OpenStudy (samigupta8):

@Michele_laino

OpenStudy (michele_laino):

is \[\huge cis\theta = \cos \theta + i\sin \theta \]

OpenStudy (samigupta8):

Yup

OpenStudy (samigupta8):

The question is like cos1°cos3°cos5°....cos89°/sin2(1)°sin2(3°)....sin2(89)°

OpenStudy (michele_laino):

then we have to evaluate this quantity: \[\huge \prod\limits_{n = 1}^{45} {\frac{{\cos \left( {2n - 1} \right)}}{{\sin \left\{ {2\left( {2n - 1} \right)} \right\}}}} \]

OpenStudy (samigupta8):

What I did was first solved the denominator term by applying sin2A =2sinAcosA And cancelled all the cos term of the denominator

OpenStudy (samigupta8):

With a degree symbol also :)

OpenStudy (michele_laino):

ok! so we have: \[\Large \prod\limits_{n = 1}^{45} {\frac{{\cos \left( {2n - 1} \right)^\circ }}{{\sin \left\{ {2\left( {2n - 1} \right)} \right\}^\circ }}} = \prod\limits_{n = 1}^{45} {\frac{1}{{2\sin \left( {2n - 1} \right)^\circ }}} \]

OpenStudy (samigupta8):

Correct!

OpenStudy (michele_laino):

I'm thinking...

OpenStudy (samigupta8):

N I m waiting..... lolxccc

OpenStudy (samigupta8):

Sir one more thing

OpenStudy (michele_laino):

ok!

OpenStudy (samigupta8):

If u luk at the terms carefully u will see that they are like 2sin1°2sin3°2sin5°...2sin85°2sin87° 2sin89° That means the last term is actually 2cos1° in disguise.. And so again if we combine the two terms first and last we get 2sin2°2sin6°....

OpenStudy (michele_laino):

I view such product in a different way, here it is: \[\begin{gathered} \prod\limits_{n = 1}^{45} {\frac{{\cos \left( {2n - 1} \right)^\circ }}{{\sin \left\{ {2\left( {2n - 1} \right)} \right\}^\circ }}} = \prod\limits_{n = 1}^{45} {\frac{1}{{2\sin \left( {2n - 1} \right)^\circ }}} = \hfill \\ \hfill \\ = \frac{1}{{{2^2}\left( {\sin 1 \cdot \sin 89} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 3 \cdot \sin 87} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 5 \cdot \sin 84} \right)}} \cdot ... \hfill \\ \hfill \\ ... \cdot \frac{1}{{2\sin 45}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. I have made a typo: \[\begin{gathered} \prod\limits_{n = 1}^{45} {\frac{{\cos \left( {2n - 1} \right)^\circ }}{{\sin \left\{ {2\left( {2n - 1} \right)} \right\}^\circ }}} = \prod\limits_{n = 1}^{45} {\frac{1}{{2\sin \left( {2n - 1} \right)^\circ }}} = \hfill \\ \hfill \\ = \frac{1}{{{2^2}\left( {\sin 1 \cdot \sin 89} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 3 \cdot \sin 87} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 5 \cdot \sin 85} \right)}} \cdot ... \hfill \\ \hfill \\ ... \cdot \frac{1}{{2\sin 45}} \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Sir, I meant the same BT forgot to write that in denominator... Now will u also do it like that only sin89°=cos1°

OpenStudy (michele_laino):

ok! I think to use this identity, for example: \[\sin 1 \cdot \sin 89 = \frac{{ - 1}}{2}\left( {\cos 90 - \cos 88} \right) = \frac{{\cos 88}}{2}\]

OpenStudy (samigupta8):

Okay!

OpenStudy (samigupta8):

What next?

OpenStudy (michele_laino):

so, I can write this: \[\begin{gathered} \prod\limits_{n = 1}^{45} {\frac{{\cos \left( {2n - 1} \right)^\circ }}{{\sin \left\{ {2\left( {2n - 1} \right)} \right\}^\circ }}} = \prod\limits_{n = 1}^{45} {\frac{1}{{2\sin \left( {2n - 1} \right)^\circ }}} = \hfill \\ \hfill \\ = \frac{1}{{{2^2}\left( {\sin 1 \cdot \sin 89} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 3 \cdot \sin 87} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 5 \cdot \sin 85} \right)}} \cdot ... \hfill \\ \hfill \\ ... \cdot \frac{1}{{2\sin 45}} = \hfill \\ \hfill \\ = \frac{1}{{2\cos 88}} \cdot \frac{1}{{2\cos 84}} \cdot \frac{1}{{2\cos 80}} \cdot ... \cdot \cdot \frac{1}{{2\cos 4}} \cdot \frac{1}{{2\sin 45}} \hfill \\ \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Yep

OpenStudy (samigupta8):

Then!

OpenStudy (michele_laino):

I'm thinking...

OpenStudy (samigupta8):

Ok

OpenStudy (michele_laino):

next step: \[\begin{gathered} \prod\limits_{n = 1}^{45} {\frac{{\cos \left( {2n - 1} \right)^\circ }}{{\sin \left\{ {2\left( {2n - 1} \right)} \right\}^\circ }}} = \prod\limits_{n = 1}^{45} {\frac{1}{{2\sin \left( {2n - 1} \right)^\circ }}} = \hfill \\ \hfill \\ = \frac{1}{{{2^2}\left( {\sin 1 \cdot \sin 89} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 3 \cdot \sin 87} \right)}} \cdot \frac{1}{{{2^2}\left( {\sin 5 \cdot \sin 85} \right)}} \cdot ... \hfill \\ \hfill \\ ... \cdot \frac{1}{{2\sin 45}} = \hfill \\ \hfill \\ = \frac{1}{{2\cos 88}} \cdot \frac{1}{{2\cos 84}} \cdot \frac{1}{{2\cos 80}} \cdot ... \cdot \cdot \frac{1}{{2\cos 4}} \cdot \frac{1}{{2\sin 45}} = \hfill \\ \hfill \\ = \frac{1}{{{2^2}\left( {\cos 4 \cdot \cos 64} \right)}} \cdot \frac{1}{{{2^2}\left( {\cos 8 \cdot \cos 68} \right)}} \cdot \frac{1}{{{2^2}\left( {\cos 12 \cdot \cos 72} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{{2^2}\left( {\cos 16 \cdot \cos 76} \right)}} \cdot \frac{1}{{{2^2}\left( {\cos 20 \cdot \cos 80} \right)}} \cdot \frac{1}{{{2^2}\left( {\cos 24 \cdot \cos 84} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{{2^2}\left( {\cos 28 \cdot \cos 88} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{{2^2}\left( {\cos 32 \cdot \cos 60} \right)}} \cdot \frac{1}{{{2^2}\left( {\cos 36 \cdot \cos 56} \right)}} \cdot \frac{1}{{{2^2}\left( {\cos 40 \cdot \cos 50} \right)}} \cdot \hfill \\ \hfill \\ \frac{1}{{{2^2}\left( {\cos 48 \cdot \cos 44} \right)}} \cdot \frac{1}{{2\sin 45}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

nevertheless it is not the right step. Here is a more useful step: \[\begin{gathered} = \frac{1}{{2\cos 88}} \cdot \frac{1}{{2\cos 84}} \cdot \frac{1}{{2\cos 80}} \cdot ... \cdot \cdot \frac{1}{{2\cos 4}} \cdot \frac{1}{{2\sin 45}} = \hfill \\ \hfill \\ = \frac{1}{{2\left( {\cos 92 + \cos 52} \right)}} \cdot \frac{1}{{2\left( {\cos 92 + \cos 44} \right)}} \cdot \frac{1}{{2\left( {\cos 92 + \cos 36} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{2\left( {\cos 92 + \cos 28} \right)}} \cdot \frac{1}{{2\left( {\cos 92 + \cos 20} \right)}} \cdot \frac{1}{{2\left( {\cos 92 + \cos 12} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{2\left( {\cos 92 + \cos 8} \right)}} \cdot \frac{1}{{2\sin 45}} = \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, we have: \[\cos 92 = \cos 90\cos 2 - \sin 90\sin 2 = - \sin 2\]

OpenStudy (michele_laino):

so, we can write: \[\begin{gathered} = - \frac{1}{{{2^7}}}\frac{1}{{\left( {\sin 2 + \cos 52} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 44} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 36} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{\left( {\sin 2 + \cos 28} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 20} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 12} \right)}} \cdot \hfill \\ \hfill \\ \cdot \frac{1}{{\left( {\sin 2 + \cos 8} \right)}} \cdot \frac{1}{{2\sin 45}} = \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. there is a typo: \[\begin{gathered} = - \frac{1}{{{2^7}}}\frac{1}{{\left( {\sin 2 + \cos 52} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 44} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 36} \right)}} \cdot \hfill \\ \cdot \frac{1}{{\left( {\sin 2 + \cos 28} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 20} \right)}} \cdot \frac{1}{{\left( {\sin 2 + \cos 12} \right)}} \cdot \hfill \\ \cdot \frac{1}{{\left( {\sin 2 + \cos 4} \right)}} \cdot \frac{1}{{2\sin 45}} \hfill \\ \end{gathered} \]

OpenStudy (samigupta8):

Sir how will we solve this now?

OpenStudy (michele_laino):

I tried to solve it this morning, nevertheless I have not any idea :(

OpenStudy (samigupta8):

Sir the options given are 1).cos1° 2).2^-1/2 3).2^1/2 4).None of these

OpenStudy (michele_laino):

ok! I will try to continue to search, and Tomorrow I will give the answer, I hope

OpenStudy (samigupta8):

Thanks sir ....

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

here is what I got: \[\huge \frac{{\sqrt 2 }}{{{2^{23}}}}\prod\limits_{n = 1}^{22} {\frac{1}{{\sin \left( {4n - 2} \right)}}} = \frac{1}{{\sqrt 2 }}\] wherein I made the last step by means of my calculator @samigupta8

OpenStudy (samigupta8):

Yes sir that'z ryt....

OpenStudy (samigupta8):

Sir without the use of calC can't we get it..?

OpenStudy (michele_laino):

yes! I think so! nevertheless I haven't been able to get the final expression